Heidelberg weekly meeting

Europe/Zurich
03.410 (Physikalisches Institut, Heidelberg)

03.410

Physikalisches Institut, Heidelberg

Description

We will have the meeting via Zoom: https://cern.zoom.us/j/92591061032?pwd=MkZzY2g1cG1YZGFwK244c210dk44dz09

Note: this is a temporary room until we get the institute account.

We will keep HeiCONF as a backup: https://heiconf.uni-heidelberg.de/sch-nqu-dd2   Access Code: 858824

Vidyo will be used as a back-up. Room extension 10823320 or use this link.

 

Tentative minute taker list at https://docs.google.com/spreadsheets/d/15oo_a_1WPxPBOtlH0LDEKG1128y1O2lfswpUAvPCQOY/edit#gid=0

Presentation by Jackob: nice ideas. Method 2 does not match method 1. Luigi: approximation of charge released in one point not valid...? Andre: diffusion might not be main contribution.

Testbeam: good pace. 5 TB of data so far. 50 Ohm*cm sensor now. At 130 V bias lot of noise. Andre: masking? Luigi: should be possible.

Masking/MP10 performance: it seems that the signal height and noise are half of Mupix8. Andre showed a plot obtained by looking at where efficiency vs threshold drops to 50% per bias voltage, and it turns at 60 V it is 120 mV, where MP8 had 230 mV. Factor of 2. Heiko: investigation needed. CSA seems ok, maybe some loss at comparator. Voltage levels might be messed. David: consider capacitance increase? Should not be effective at 200 Ohm*cm.

Alena: nothing new.

Daniel: finished pin table of FEB. Finished first part of script, which takes table and produces file for pin assignment and top level entity. Next week check syntax errors, bugs,... and then integrated generated files in quartus.

Heiko: more feedback from Optim. Next Friday they will discuss it internally. CMP will leave residuals which we would have to clean. They are pessimistic. Luigi: investigated with DISCO, they offer Chemical-free stress relief dry polishing. Andre: should be same as Alice. Luigi: should try. Jan: last experience with DISCO was bad, all wafers broken. Heiko: could try, then check IV curves. Main point does not change, we cannot deplete more than 30 V anyway.

Jackob: no more than presentation.

Jan: Mupix10 unirradiated does not work. Something must be wrong. After testbeam will get in touch to check. Atlaspix3 works. Cooling system does not look to work very well. Ordered atlaspix3 inserts.

Sebastian: not much. Check in the lab to reproduce Zak issue, but looks not reproducable. Waiting for people back from testbeam.

Stephan: nothing new. Thesis writing.

Thomas: integration run analysis: looking for MuSR signal. Maybe Marius found something. We replaced the target with something magnetized, and check precession. As for Mu3e, in principle it is possible for data with magnet on, but needs more data polishing. Si-heater: previous measurement are invalid, there is a parasitic resistance that affects the temperature measurements. Now measurements good.

There are minutes attached to this event. Show them.
    • 2:00 PM 2:20 PM
      Presentations
      Convener: Luigi Vigani (University of Heidelberg)
    • 2:20 PM 3:05 PM
      Round Table 45m

      Reports of work progress

      Speakers: Adrian Herkert (Physikalisches Institut, Uni Heidelberg) , Adrian Stoll, Alena Weber (Universität Heidelberg) , Andre Schoening (University Heidelberg, Institute of Physics) , Annie Meneses Gonzalez (PhD Student) , Benjamin Weinläder (PI Heidelberg) , David Maximilian Immig (Universität Heidelberg) , Dohun Kim (Physikalisches Institut University of Heidelberg) , Florian Frauen, Frank Meier Aeschbacher (PSI - Paul Scherrer Institut) , Heiko Augustin (PI Uni Heidelberg) , Jakob Stricker, Lars Noehte (PI Heidelberg) , Luigi Vigani (University of Heidelberg) , Lukas Mandok (Physikalisches Institut Heidelberg) , Marius Menzel (Institute of Physics, University of Heidelberg) , Sebastian Dittmeier (Physikalisches Institut, Universität Heidelberg) , Sebastian Preuß (PI Uni Heidelberg) , Stephan Lachnit, Thomas Theodor Rudzki (Physikalisches Institut Heidelberg)
    • 3:05 PM 3:15 PM
      AOB